# Retrieval Augmented Generation for Dynamic Graph Modeling

### v1

@@ 4.2 Graph Fusion-based Generator
$e_{\text{fus}}$ $=\mathtt{MeanPooling}(\text{GCN}(G_{\text{fus}})),$ (9)
$\tilde{x}_{q}$ $=[e_{\text{fus}}\parallel x_{q}],$ (10)
@@
Training and inference.
We adopt the same sequence model with the same training objective (Wu, Fang, and Liao 2024) as in Fig. 2(a). During training, we freeze the parameters of the sequence model, except for the output layer which is updated along with the GCN parameters used for graph fusion.
During testing, we first apply the retriever model to retrieve top-K demonstrations for each query as introduced in Sect. 4.1. Then we perform graph fusion on these demonstrations and concatenate the fused graph representation with the query sequence as illustrated in Eqs. (9) and (10). The concatenated sequence is subsequently fed into the trained sequence model for link prediction.
5 Experiment
@@
5.1 Experimental Setup
Datasets.
We evaluate the performance of the proposed model on three datasets from different domains including the communication network UCI (Panzarasa, Opsahl, and Carley 2009), the citation network Hepth (Leskovec, Kleinberg, and Faloutsos 2005), and the multi-turn task-oriented conversation dataset MMConv (Liao et al. 2021). We follow the same dataset preprocessing as SimpleDyG (Wu, Fang, and Liao 2024). It is important to note that the ML-10M dataset was not used because it is not suitable for RAG; we observed no performance enhancement even when the query was augmented with ground-truth demonstrations. We summarize the statistics of the three datasets in Table 1.
Dataset UCI Hepth MMConv
Domain Social Citation Conversation
# Nodes 1,781 4,737 7,415
# Edges 16,743 14,831 91,986
Table 1: Dataset statistics.
Method UCI Hepth MMConv
Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
DySAT 0.009±0.003 0.010±0.003 0.010±0.001 0.008±0.004 0.007±0.002 0.005±0.001 0.108±0.089 0.102±0.085 0.095±0.080
EvolveGCN 0.072±0.046 0.064±0.045 0.032±0.026 0.008±0.002 0.009±0.004 0.007±0.002 0.050±0.015 0.051±0.021 0.032±0.017
DyRep 0.009±0.008 0.011±0.018 0.010±0.005 0.009±0.006 0.031±0.024 0.010±0.006 0.156±0.054 0.140±0.057 0.067±0.025
JODIE 0.018±0.019 0.022±0.023 0.012±0.009 0.010±0.008 0.031±0.021 0.011±0.008 0.052±0.039 0.041±0.016 0.032±0.022
TGAT 0.022±0.004 0.061±0.007 0.020±0.002 0.011±0.007 0.034±0.023 0.011±0.006 0.118±0.004 0.089±0.033 0.058±0.021
TGN 0.014±0.010 0.041±0.017 0.011±0.003 0.011±0.006 0.030±0.012 0.008±0.001 0.085±0.050 0.096±0.068 0.066±0.038
TREND 0.083±0.015 0.067±0.010 0.039±0.020 0.010±0.008 0.031±0.003 0.010±0.002 0.134±0.030 0.116±0.020 0.060±0.018
GraphMixer 0.097±0.019 0.104± 0.013 0.042 ±0.005 0.009±0.002 0.011±0.008 0.010±0.003 0.206±0.001 0.172±0.029 0.085±0.016
SimpleDyG 0.109±0.014 0.104±0.010 0.092±0.014 0.013±0.006 0.035±0.014 0.013±0.006 0.170±0.010 0.184±0.012 0.169±0.010
RAG4DyG 0.111±0.013 0.122±0.014 0.097±0.010 0.019±0.002 0.045±0.003 0.019±0.002 0.194±0.005 0.208±0.005 0.194±0.005
@@ 5.1 Experimental Setup
Table 2: Performance comparison for dynamic link prediction. (Best results are bolded; runners-up are underlined.)
Baselines.
@@
We compare RAG4DyG with the state-of-the-art approaches of dynamic graph modeling. (1) discrete-time approaches: DySAT (Sankar et al. 2020) and EvolveGCN (Pareja et al. 2020); (2) continuous-time approaches: DyRep (Trivedi et al. 2019), JODIE (Kumar, Zhang, and Leskovec 2019), TGAT (Xu et al. 2020), TGN (Rossi et al. 2020), TREND (Wen and Fang 2022), GraphMixer (Cong et al. 2022) and SimpleDyG (Wu, Fang, and Liao 2024). The details of the baselines can be found in Appendix B.
Implementation Details.
@@
Following the method outlined in (Cong et al. 2022; Wu, Fang, and Liao 2024), we represent the dynamic graph as an undirected graph. We split all datasets into training, validation, and test sets based on temporal sequence same as SimpleDyG (Wu, Fang, and Liao 2024). Given ${T}$ timesteps in each dataset, the data at the final timestep ${T}$ is used as the testing set, the data at ${T-1}$ is served as the validation set, and the remaining data from earlier timesteps is used for training. All training data including the retrieval pool for the retriever and generator is drawn from this training data split. For retrieval augmented generation model training, we first train SimpleDyG without augmentation using the reported parameter in this paper. Then we fix the parameters of SimipleDyG except for the last linear layer and fine-tune them with the GCN model. The number of GCN layers in the generator model is 1 for all datasets. We repeat each experiment 10 times and report the average results along with the standard deviation. The number of demonstrations is 7 for all datasets. Hyper-parameter tuning is performed on the validation set and we provide the settings in Appendix D.
Evaluation Metrics.
@@
We evaluate the performance upon three metrics Recall@5, NDCG@5 and Jaccard (Wu, Fang, and Liao 2024). Recall@5 and NDCG@5 are widely used in ranking tasks measuring the quality of top-ranked predictions (Wang et al. 2019). Recall@5 assesses the proportion of relevant nodes in the top 5 predictions, while NDCG@5 evaluates the ranking quality by considering the positions of the relevant nodes. Jaccard (Jaccard 1901) measures the similarity between the predicted and ground truth sequences by comparing the intersection and union of the two sets.
5.2 Performance Comparison
@@
We evaluate the performance of RAG4DyG for the dynamic link prediction task, and the results compared to the state-of-the-art baselines are presented in Table 2. We make the following observations.
@@
The proposed RAG4DyG generally outperforms all baselines across different datasets under the three metrics. In particular, compared to SimpleDyG, which is also our backbone, RAG4DyG consistently shows superior performance, highlighting the effectiveness of our retrieval-augmented generation framework. Note that GraphMixer performs slightly better in Recall@5 on the MMConv dataset, but its significantly lower performance in NDCG@5 and Jaccard indicates that its predictions are not ranked optimally or maintaining the overall set integrity compared to RAG4DyG. This indicates that RAG4DyG can better model the temporal and contextual relationships due to the specific design of the retriever and generator.
Figure 3: Ablation study for retrieval results.
@@
Second, RAG4DyG exhibits significant advantages in inductive scenarios such as the Hepth dataset. This setting is particularly challenging because it involves nodes not appearing during the training phase, requiring the model to generalize to entirely new structures and relationships. RAG4DyG’s success is attributed to its retrieval-augmented mechanism, which enhances the model’s ability to generalize by providing rich contextual information relevant to the new, unseen nodes. Unlike models that rely solely on the immediate neighborhood or predefined structures, RAG4DyG dynamically adapts to the new nodes, ensuring that the predictions are guided by the most relevant and similar historical data.
5.3 Model Analysis
@@
Ablation Study.
To evaluate the effectiveness of different modules in the retrieval model, we compare RAG4DyG with two variants w/o CCL and w/o Decay which exclude the context-aware contrastive learning and time decay component in the retrieval model. We evaluate the performance for both retrieval and link prediction tasks. We use HR@k (Hit Ratio@k) metrics for the retrieval model, measuring the proportion of cases where at least one of the top-k retrieved items is relevant. As shown in Fig. 3 and 4, the full model outperforms the two variants, underscoring the benefits of incorporating context-aware contrastive learning and time decay modulation.
Figure 4: Ablation study for link prediction results.
Effect of Different Retrieval Methods.
To further investigate the effectiveness of the retrieval model, we compare our model with two different retrieval methods, namely, BM25 and Jaccard, in Table 3 and 4. In Table 4, we only report NDCG@5 and present the remaining metrics in Appendix C. BM25 (Fang, Tao, and Zhai 2004) is an extension of the Term Frequency-Inverse Document Frequency (TF-IDF) model, which calculates a relevance score between the query sequence and each candidate sequence in the retrieval pool. The relevance score is derived from the occurrence frequency of the nodes in the query and the retrieval pool. Jaccard (Jaccard 1901) measures the similarity between two sets by comparing the size of their intersection to the size of their union. Note that in the citation dataset Hepth, all queries in the test set only contain unseen target nodes that never appear in the retrieval pool and have no historical interactions. As a result, the BM25 and Jaccard scores between the queries and the candidates in the retrieval pool are always zeros. On the other hand, our retrieval model is trained based on the sequence representations. For a query sequence containing only the target node, we can still obtain its representation using the sequence model trained for the retrieval model, and further calculate its contextual similarity with the candidate sequences in the retrieval pool.
Method UCI Hepth MMConv
HR@1 HR@3 HR@1 HR@3 HR@1 HR@3
BM25 0.100 0.136 - - 0.025 0.233
Jaccard 0.100 0.109 - - 0.007 0.213
RAG4DyG 0.096 0.153 0.025 0.049 0.034 0.168
“-” denotes that the method is unable to perform retrieval. The reason is explained in the corresponding description of this table in Sect. 5.3.
Table 3: Retrieval performance of various retrieval methods.
Method UCI Hepth MMConv
BM25 0.121±0.009 - 0.207±0.003
Jaccard 0.$1$ 0.207±0.005
RAG4DyG 0.122±0.014 0.045±0.003 0.208±0.005
GroundTruth 0.129±0.010 0.062±0.007 0.224±0.008
@@
Table 4: Generative accuracy of various retrieval methods.
In Table 3, we analyze the retrieval performance of different retrieval methods. Our retrieval model shows better performance than other retrieval strategies. Notably, in inductive scenarios like the Hepth dataset, BM25 and Jaccard fail to work with new query nodes lacking historical interactions. In contrast, our model can handle them effectively and achieve solid performance. The high $HR@3$ performance of BM25 and Jaccard on the MMConv dataset can be attributed to the nature of the dialogue dataset, where temporal order is less critical, and certain nodes associated with specific slot values are more discriminative.
@@
Effect of the Number of Demonstrations K.

### v2

@@ 4.2. Graph Fusion-based Generator
(9) $e_{\text{fus}}$ $=\mathtt{MeanPooling}(\text{GCN}(G_{\text{fus}})),$
(10) $\tilde{x}_{q}$ $=[e_{\text{fus}}\parallel x_{q}],$
@@
Training and Inference. We adopt the same sequence model with the same training objective (Wu et al., 2024) as in Fig. 2(a). During training, we freeze the parameters of the sequence model, except for the output layer which is updated along with the GCN parameters used for graph fusion. During testing, we first apply the retriever model to retrieve top-K demonstrations for each query as introduced in Sec. 4.1. Then we perform graph fusion on these demonstrations and concatenate the fused graph representation with the query sequence as illustrated in Eq. (9) and (10). The concatenated sequence is subsequently fed into the trained model for link prediction.
5. Experiment
@@
5.1. Experimental Setup
5.1.1. Datasets.
We evaluate the performance of the proposed model on six datasets from different domains, including a communication network UCI (Panzarasa et al., 2009), a citation network Hepth (Leskovec et al., 2005), a multi-turn task-oriented conversation dataset MMConv (Liao et al., 2021), a behavioral interaction network Wikipedia (Huang et al., 2024), an email network Enron (Zhang et al., 2024), and a hyperlink network Reddit (Kumar et al., 2018). We summarize the statistics of these datasets in Table 1. We follow the preprocessing steps of SimpleDyG to map the dynamic graphs into sequences with special tokens (Wu et al., 2024). Notably, the Hepth and Reddit datasets exhibit an inductive nature, as they contain previously unseen nodes with no historical interactions. The details of the Wikipedia, Enron and Reddit datasets are provided below, while information on the remaining datasets can be found in SimpleDyG (Wu et al., 2024).
  • Wikipedia (Huang et al., 2024): This dataset captures the co-editing activity on Wikipedia pages over one month. It is a bipartite interaction network in which editors and wiki pages serve as nodes. Each edge corresponds to an interaction where a user edits a page at a specific timestamp. To facilitate temporal sequence alignment, the dataset is divided into 16 time steps based on the timestamps of the interactions.
  • Enron (Zhang et al., 2024): This dataset represents the email communications among employees of Enron Corporation over three years (1999–2002). Nodes represent employees, while edges correspond to emails exchanged between them ordered by the sending timestamps of the emails. For temporal sequence alignment, we split the dataset into 17 time steps based on the timestamps.
  • Reddit (Kumar et al., 2018): This dataset represents a subreddit-to-subreddit hyperlink network, derived from timestamped posts containing hyperlinks between subreddits. We focus on hyperlink data within the body of posts, covering the period from 2016 to 2017. The dataset is divided into 12 time steps for temporal sequence alignment based on the post timestamps.
Table 1. Dataset statistics.
Dataset UCI Hepth MMConv Wikipedia Enron Reddit
Domain Social Citation Conversation Behavior Social Hyperlink
# Nodes 1,781 4,737 7,415 9,227 42,711 11,901
# Edges 16,743 14,831 91,986 157,474 797,907 62,919
@@ 5.1. Experimental Setup
Table 2. Performance comparison for dynamic link prediction with mean and standard deviation across 10 runs. Best results are bolded; runners-up are underlined. * indicates that our model significantly outperforms the best baseline based on the two-tail $t$-test ($p<0.05$).
Datasets Models DySAT EvolveGCN DyRep JODIE TGAT TGN TREND GraphMixer IDOL SimpleDyG RAG4DyG
UCI Recall@5 0.009±0.003 0.072±0.046 0.009±0.008 0.018±0.019 0.022±0.004 0.014±0.010 0.083±0.015 0.097±0.019 0.093±0.029 0.109±0.014 0.111±0.013
NDCG@5 0.010±0.003 0.064±0.045 0.011±0.018 0.022±0.023 0.061±0.007 0.041±0.017 0.067±0.010 0.104±0.013 0.075±0.022 0.104±0.010 0.122*±0.014
Jaccard 0.010±0.001 0.032±0.026 0.010±0.005 0.012±0.009 0.020±0.002 0.011±0.003 0.039±0.020 0.042±0.005 0.014±0.002 0.092±0.014 0.097±0.010
Hepth Recall@5 0.008±0.004 0.008±0.002 0.009±0.006 0.010±0.008 0.011±0.007 0.011±0.006 0.010±0.008 0.009±0.002 0.007±0.002 0.013±0.006 0.019*±0.002
NDCG@5 0.007±0.002 0.009±0.004 0.031±0.024 0.031±0.021 0.034±0.023 0.030±0.012 0.031±0.003 0.011±0.008 0.011±0.003 0.035±0.014 0.045*±0.003
Jaccard 0.005±0.001 0.007±0.002 0.010±0.006 0.011±0.008 0.011±0.006 0.008±0.001 0.010±0.002 0.010±0.003 0.006±0.001 0.013±0.006 0.019*±0.002
MMConv Recall@5 0.108±0.089 0.050±0.015 0.156±0.054 0.052±0.039 0.118±0.004 0.085±0.050 0.134±0.030 0.206±0.001 0.169±0.006 0.170±0.010 0.194±0.005
NDCG@5 0.102±0.085 0.051±0.021 0.140±0.057 0.041±0.016 0.089±0.033 0.096±0.068 0.116±0.020 0.172±0.029 0.115±0.039 0.184±0.012 0.208*±0.005
Jaccard 0.095±0.080 0.032±0.017 0.067±0.025 0.032±0.022 0.058±0.021 0.066±0.038 0.060±0.018 0.085±0.016 0.015±0.002 0.169±0.010 0.194*±0.005
Wikipedia Recall@5 0.003±0.005 0.012±0.01 0.003±0.002 0.017±0.005 0.006±0.004 0.016±0.018 0.022±0.012 0.010±0.007 0.022±0.008 0.356±0.016 0.369*±0.006
NDCG@5 0.002±0.003 0.008±0.007 0.002±0.002 0.015±0.003 0.005±0.005 0.015±0.022 0.016±0.018 0.007±0.006 0.015±0.005 0.398±0.03 0.389±0.008
Jaccard 0.001±0.001 0.004±0.004 0.001±0.001 0.007±0.002 0.002±0.002 0.007±0.009 0.007±0.021 0.004±0.002 0.004±0.001 0.320±0.027 0.328±0.007
Enron Recall@5 0.002±0.004 0.004±0.011 0.021±0.001 0.005±0.005 0.020±0.002 0.001±0.001 0.023±0.003 0.021±0.002 0.024±0.014 0.094±0.005 0.100*±0.003
NDCG@5 0.001±0.002 0.007±0.020 0.036±0.002 0.061±0.039 0.036±0.001 0.003±0.001 0.027±0.001 0.037±0.001 0.025±0.011 0.114±0.005 0.119*±0.004
Jaccard 0.001±0.001 0.003±0.009 0.019±0.001 0.011±0.007 0.020±0.001 0.001±0.001 0.012±0.001 0.020±0.002 0.008±0.003 0.068±0.003 0.071*±0.002
Reddit Recall@5 0.001±0.002 0.006±0.002 0.019±0.004 0.013±0.003 0.001±0.001 0.001±0.001 0.002±0.003 0.001±0.001 0.003±0.002 0.101±0.019 0.119*±0.006
NDCG@5 0.001±0.002 0.012±0.003 0.020±0.004 0.015±0.002 0.001±0.001 0.002±0.001 0.003±0.002 0.003±0.001 0.005±0.003 0.134±0.012 0.143±0.005
Jaccard 0.001±0.001 0.003±0.001 0.013±0.004 0.007±0.002 0.001±0.001 0.001±0.002 0.001±0.001 0.002±0.001 0.002±0.001 0.088±0.012 0.096±0.003
5.1.2. Baselines.
@@
We compare our model RAG4DyG with the state-of-the-art dynamic graph models, which include (1) discrete-time approaches: DySAT (Sankar et al., 2020) and EvolveGCN (Pareja et al., 2020); (2) continuous-time approaches: DyRep (Trivedi et al., 2019), JODIE (Kumar et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), TREND (Wen and Fang, 2022), GraphMixer (Cong et al., 2022), IDOL (Zhu et al., 2024) and SimpleDyG (Wu et al., 2024).
  • DySAT (Sankar et al., 2020) utilizes self-attention mechanisms to capture both structural and temporal patterns in dynamic graphs through discrete-time snapshots.
  • EvolveGCN (Pareja et al., 2020) leverages recurrent neural networks to model the evolution of the parameters of a graph convolutional network over discrete time steps.
  • DyRep (Trivedi et al., 2019) models dynamic graphs in continuous time by incorporating both temporal point processes and structural dynamics to capture interactions and node dynamics.
  • JODIE (Kumar et al., 2019) focuses on user and item embedding trajectories over continuous time, predicting future interactions by modeling user and item embeddings jointly.
  • TGAT (Xu et al., 2020) employs temporal graph attention layers and time encoding to capture temporal dependencies and structural information for dynamic graphs.
  • TGN (Rossi et al., 2020) combines GNNs with memory modules to maintain node states over continuous time, effectively learning from dynamic interactions.
  • TREND (Wen and Fang, 2022) integrates temporal dependencies based on the Hawkes process and GNNs to learn the dynamics of graphs.
  • GraphMixer (Cong et al., 2022) introduces a novel architecture that leverages MLP-mixer to learn link-encoder and node encoder for evolving graphs in continuous time.
  • IDOL (Zhu et al., 2024) is a contrastive learning-based model tailored for dynamic graph representation learning. It utilizes a Personalized PageRank-based algorithm to incrementally update the node embedding and adopt a topology-monitorable sampling method to generate contrastive pairs for efficient training.
  • SimpleDyG (Wu et al., 2024) reformulates the dynamic graph modeling as a sequence modeling task and mapped the dynamic interactions of target nodes as sequences with specially designed tokens. It simplifies dynamic graph modeling without complex architectural changes to effectively capture temporal dynamics.
5.1.3. Implementation Details.
@@
Following the method outlined in (Cong et al., 2022; Wu et al., 2024), we represent the dynamic graph as an undirected graph. We split all datasets into training, validation, and test sets based on temporal sequence same as SimpleDyG (Wu et al., 2024). Given ${T}$ timesteps in each dataset, the data at the final timestep ${T}$ is used as the testing set, the data at ${T-1}$ is served as the validation set, and the remaining data from earlier timesteps is used for training. All training data including the retrieval pool for the retriever and generator is drawn from this training data split. For retrieval augmented generation model training, we first train SimpleDyG without augmentation using the finetuned parameters. Then we fix the parameters of SimipleDyG except for the last linear layer and fine-tune them with the GCN model. The number of GCN layers in the generator model is 1 for all datasets. We repeat each experiment 10 times and report the average results along with the standard deviation. The number of demonstrations is 7 for all datasets.
To facilitate retrieval model training, we regard the samples in the training dataset as our retrieval pool $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ where each pair $(x_{i},y_{i})$ represents the historical sequence and its corresponding target sequence. Specifically, $x_{i}$ is the input sequence before the last time step and $y_{i}$ is the output sequence at the last time step. We annotate demonstrations based on the Jaccard similarity between the output sequences among all the pairs in $D$. The Jaccard similarity of two output sequences, $y_{i}$ and $y_{j}$, is given by
(11) $r(y_{i},y_{j})=\frac{|y_{i}\cap y_{j}|}{|y_{i}\cup y_{j}|}.$
To control the quality of annotated data, we set a threshold of 0.8 to select highly similar demonstrations for each sample. These filtered annotations are then used to train the retriever model. The number of training samples for the UCI, Hepth, MMConv, Wikipedia, Enron and Reddit datasets are 9 578, 8 250, 10 762, 162 408, 2 510 666 and 185 764, respectively.
We ran all the experiments on a Nvidia L40 GPU and tuned the hyper-parameters for all the methods based on the validation set. For all the baselines, we tuned the models based on the hyper-parameters reported in their papers. For our RAG4DyG method, we set the number of layers, heads and dimensions of hidden states for the backbone SimpleDyG to $(6,8,768)$, $(12,2,256)$, $(2,2,256)$, $(2,6,768)$, $(2,6,768)$, and $(2,8,512)$ across the six datasets. The time decay rate $\lambda$ for the retrieval model in Eq. (4) was tuned according to the time granularity of different datasets, with days for the UCI dataset, months for the Hepth dataset, turns for the MMConv dataset, and hours for other datasets. We explored a range of values $\lambda=\{10^{-4},10^{-3},10^{-2},10^{-1},1,10,100\}$, ultimately selecting $\lambda=10^{-4}$ for UCI, $\lambda=0.1$ for Hepth, $\lambda=10$ for MMConv and Enron, and $\lambda=1$ for Wikipedia and Reddit datasets. The coefficient $\alpha$ in the loss function in Eq. (8) was tuned across $\{0.2,0.4,0.6,0.8,1\}$, resulting in final values of $\alpha=1$ for UCI and MMConv, $\alpha=0.4$ for Hepth, and $\alpha=0.2$ for Wikipedia, Enron and Reddit. Additional parameter settings for the three datasets are as follows: the temperature $\tau$ in the two contrastive learning tasks for all datasets was set to $\tau=0.1$, the batch size of the retriever model for all datasets was set to $N=128$, and the masking and cropping portions in context-aware contrastive learning were set to 0.8, 0.8, 0.8, 0.6, 0.6, 0.2 and 0.4, 0.6, 0.6, 0.8, 0.8, 0.8 across the six datasets, respectively.
5.1.4. Evaluation Metrics.
@@
Inspired by SimpleDyG (Wu et al., 2024), we assess the performance of our approach and baselines using three key metrics: Recall@5, NDCG@5, and Jaccard (Wu et al., 2024). Recall@5 and NDCG@5 are commonly employed in ranking tasks to evaluate the quality of top-ranked predictions (Wang et al., 2019). Specifically, Recall@5 measures the proportion of relevant nodes that appear among the top five predictions, while NDCG@5 considers the ranking positions of the relevant nodes to provide a more nuanced assessment of ranking quality. Additionally, the Jaccard index (Jaccard, 1901) quantifies the similarity between the predicted and ground truth sequences by calculating the ratio of their intersection to their union.
Figure 3. Ablation study for retrieval results.
Figure 4. Ablation study for link prediction results.
5.2. Performance Comparison
@@
We assess the performance of RAG4DyG on the dynamic link prediction task, with the results benchmarked against state-of-the-art baselines, as shown in Table 2. Our analysis reveals the following key observations.
@@
First of all, the proposed RAG4DyG generally outperforms all baselines across different datasets under the three metrics. In particular, compared to SimpleDyG, which is also our backbone, RAG4DyG consistently shows superior performance, highlighting the effectiveness of our retrieval-augmented generation framework. Note that GraphMixer performs slightly better in Recall@5 on the MMConv dataset, but its significantly lower performance in NDCG@5 and Jaccard indicates that its predictions are not ranked optimally or maintaining the overall set integrity compared to RAG4DyG. This indicates that RAG4DyG can better model the temporal and contextual relationships due to the specific design of the retriever and generator. Generally speaking, the performance of SimpleDyG and RAG4DyG which reformulate the dynamic graph link prediction as a sequence generation task show promising performance compared with node pair ranking-based baselines, especially on the Wikipedia dataset which contains a higher frequency of repeated interaction behaviors. This characteristic makes sequence-based models more effective, as they can leverage the temporal consistency and recurrent patterns in the data to better capture the underlying dynamics of the graph.
@@
Second, RAG4DyG exhibits significant advantages in inductive scenarios such as the Hepth and Reddit datasets. This setting is particularly challenging because it involves nodes not seen during training, requiring the model to generalize to entirely new structures and relationships. RAG4DyG’s success is attributed to its retrieval-augmented mechanism, which enhances the model’s ability to generalize by providing rich contextual information relevant to the new, unseen nodes. Unlike models that rely solely on the immediate neighborhood or predefined structures, RAG4DyG dynamically adapts to the new nodes, ensuring that the predictions are guided by the most relevant and similar historical data.
5.3. Model Analysis
@@
Table 3. Retrieval performance of various retrieval methods.
Method UCI Wikipedia Hepth Reddit
HR@1 HR@3 HR@7 HR@1 HR@3 HR@7 HR@1 HR@3 HR@7 HR@1 HR@3 HR@7
BM25 0.100 0.136 0.200 0.369 0.405 0.488 - - - - - -
Jaccard 0.100 0.109 0.146 0.369 0.445 0.430 - - - - - -
RAG4DyG 0.100 0.155 0.164 0.369 0.455 0.523 0.025 0.049 0.073 0.180 0.218 0.228
“-” denotes that the method is unable to perform retrieval. The reason is explained in the corresponding description of this table in Sec. 5.3.
Table 4. Generative performance of various retrieval methods.
Method UCI Wikipedia Hepth Reddit
Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
BM25 0.111±0.007 0.121±0.009 0.093±0.004 0.368±0.01 0.389±0.012 0.325±0.01 - - - - - -
Jaccard 0.104±0.009 0.113±0.011 0.088±0.010 0.368±0.013 0.388±0.014 0.321±0.011 - - - - - -
RAG4DyG 0.111±0.013 0.122±0.014 0.097±0.010 0.369±0.006 0.389±0.008 0.328±0.007 0.019±0.002 0.045±0.003 0.019±0.002 0.119±0.006 0.143±0.005 0.096±0.003
GroundTruth 0.121±0.010 0.129±0.010 0.107±0.012 0.390±0.008 0.400±0.007 0.340±0.006 0.028±0.004 0.062±0.007 0.028±0.004 0.121±0.008 0.145±0.008 0.099±0.005
@@
5.3.1. Ablation Study.
To evaluate the effectiveness of different modules in the retrieval model, we compare RAG4DyG with two variants w/o CCL and w/o Decay which exclude the context-aware contrastive learning and time decay component in the retrieval model. We evaluate the performance for both retrieval and link prediction tasks. We use HR@k (Hit Ratio@k) metrics for the retrieval model, measuring the proportion of cases where at least one of the top-k retrieved items is relevant. As shown in Fig. 3 and 4, the full model outperforms the two variants, underscoring the benefits of incorporating context-aware contrastive learning and time decay modulation. Notably, the w/o Decay variant exhibits the worst performance across both tasks, emphasizing the critical role of time decay in capturing temporal relevance and accurately modeling the evolving dynamics of the graph.
5.3.2. Effect of Different Retrieval Methods.
To further investigate the effectiveness of the retrieval model, we compare our model with two different retrieval methods, namely, BM25 and Jaccard, in Table 3 and 4. BM25 (Fang et al., 2004) is an extension of the Term Frequency-Inverse Document Frequency (TF-IDF) model, which calculates a relevance score between the query sequence and each candidate sequence in the retrieval pool. The relevance score is derived from the occurrence frequency of the nodes in the query and the retrieval pool. Jaccard (Jaccard, 1901) measures the similarity between two sets by comparing the size of their intersection to the size of their union. Note that in the citation dataset Hepth and hyperlink dataset Reddit, the queries in the test set contain unseen target nodes that never appear in the retrieval pool and have no historical interactions. As a result, the BM25 and Jaccard scores between the queries and the candidates in the retrieval pool are always zeros. On the other hand, our retrieval model is trained based on the sequence representations. For a query sequence containing only the target node, we can still obtain its representation using the sequence model trained for the retrieval model, and further calculate its contextual similarity with the candidate sequences in the retrieval pool.
Figure 5. Effect of the number of demonstrations K.
In Table 3, we analyze the retrieval performance of different methods. For transductive scenarios, our retrieval model shows comparable performance to other retrieval strategies. Notably, in inductive scenarios like the Hepth and Reddit datasets, BM25 and Jaccard fail to work with new query nodes lacking historical interactions. In contrast, our model can handle them effectively and achieve solid performance.
@@
5.3.3. Effect of the Number of Demonstrations K